# What the flock knows that the birds do not: exploring the emergence of joint agency in multi-agent active inference

### v1

@@ 5 Discussion
In the first analysis, we used the concept of (nested) Markov blankets to formalize how agency can extend beyond individual active inference agents to groups of agents—that is, from birds to flocks. We showed that during flocking, the Markov blanket, which formalizes a statistical boundary around an agent, can expand to enclose a set of birds forming a collective unit. Within the flock, subsets of birds play the roles of internal, sensory, and active states of a collective Markov blanket, while others constitute external states. This analysis extends prior work on collective, multi-agent systems—including groups of neurons (palacios2019emergence), collections of primordial cells (friston2013life), and developing tissues (friston2015knowing)—by showing that the notion of the Markov blanket can account for situations in which each agent retains its individual autonomy, while the collective simultaneously acquires a higher-level, joint agency. From this perspective, agency is not a fixed attribute of individual systems but an emergent property of coupled dynamics that maintain conditional independence from their surroundings. As interactions among agents strengthen, the boundaries that define “self” can expand or contract, giving rise to transient or stable forms of joint agency. This provides a unified statistical and dynamical framework for understanding how perception, action, and decision-making scale from individuals to collectives—from neurons forming neural assemblies to organisms forming social or ecological systems.
@@
The simulations revealed that larger flocks reacted more rapidly to the predator, as indicated by a faster onset of destabilization during later stages, even though the rules governing the propagation of “stress” did not depend on the level of organization.
@@
Most importantly, our results show that the flock as a whole possesses synergistic information about the predator’s location, and that this information increases with flock size and extends farther from the predator. This suggests a formally grounded—albeit implicit—notion of collective knowledge (or shared world model) that goes beyond the information available to individual birds.
@@
Summing up, our simulations illustrate a simple yet general method to characterize a collective agent, such as a flock, as an entity that maintains statistical separation from its environment through a Markov blanket, and exhibits knowledge beyond its components through synergistic information. More broadly, these results illustrate how collective entities can extend the spatial and temporal reach of perception, cognition, and action—or their cognitive light cone (levin2019computational)—beyond the boundaries of any single agent.
@@
There is, however, a conceptual distinction between the implicit form of collective knowledge revealed by synergistic information and the explicit forms typically studied in cognitive science. In cognitive and social sciences, collective knowledge is often assumed to rely on explicit internal models of others’ mental states or shared epistemic representations (e.g., “common ground” or “what we all know that we know”). Similarly, our simulations demonstrate collective sensing and action that do not depend on such explicit representations but instead arise from simple interaction rules. In human societies, collective action is often defined by intentionality—individuals act with the goal of producing collective outcomes. By contrast, in simpler biological systems, such as morphogenesis or pattern formation, intentionality can be replaced by teleological organization: systems are structured in such a way to produce collective outcomes even without explicit goals.
@@
An open question, then, is how to bridge our minimal notion of joint agency with more sophisticated accounts used in cognitive science. A natural starting point is the good regulator theorem (conant1970every), which states that any effective controller must embody a model of the system it regulates. The multi-agent systems we described do not have a model of their collective behavior; rather, they collectively are or become a model of social action through their pattern of couplings. More complex agents, by contrast, possess explicit internal models of collective behavior, with shared representations, goals, and intentions. Understanding how living systems evolve from being to having a model may shed light on the developmental and evolutionary origins of collective cognition.
@@
Finally, this perspective invites consideration of agency at even broader spatial and temporal scales, such as in niche construction (pio2025scale). Extending these ideas across generations and cultural dynamics raises deeper questions (pezzulo2025shared; clark1998extended; Friston2024): To what extent are our minds shaped by social dynamics? How far do our Markov blankets extend to encompass others? Are our generative models so profoundly influenced by social and cultural processes that purely individual agency becomes impossible to isolate? And does the fact that knowledge is distributed across people, artifacts, and institutions—books, tools, and the internet—render minds inherently socially extended and embedded by design? We leave these questions open for future research.
@@ A.1 Active inference framework
In this work, we consider birds as Bayesian agents with sensory, active, and internal states that are updated according to the Active Inference principles (Parr2022). Active inference (AIF) is a theoretical framework that provides a unified account of perception, action, and learning of both living and artificial systems. It posits that an agent’s behavior can be understood as the process of maximizing the evidence for the implicit statistical model of the world it embodies, in response to the sensory information flow, by selecting adaptive sequences of actions.
@@
We adopted a hybrid modeling approach. We borrowed the Potts model (wu1982potts), a statistical physics model describing spin glasses in which spins are arranged on a lattice, a periodic graph, and combined it with the computational model developed by Reynolds to mimic collective behaviors in computer vision applications (reynolds1987flocks).
The Potts model is a generalization of the Ising model (ising1925) for $q>2$ components. It has been used to address numerous problems in collective behavior, which are often described as instances of lattice statistics. Indeed, replacing continuous symmetry with discrete symmetry allows for a simpler, more tractable understanding of the flocking transition. Recently, Solon and Tailleur argued that flocking can be described as a transition from a disordered phase to a polar-ordered phase in an active Ising model, in which spins both diffuse and align on the lattice, a coarse-grained representation of the space (solon2013revisiting). (chatterjee2020flocking) studied a square lattice in which active particles have four internal states corresponding to the four directions of motion. A local alignment rule inspired by the ferromagnetic 4-state Potts model, combined with self-propulsion via biased diffusion based on the internal particle states, leads to flocking at high densities and low noise.
@@ A.2 Generative model for flocking
    where $\theta_{ij}$ is the angle associated to the two-dimensional unit position vector $(\mathbf{r}_{i}-\mathbf{r}_{j})=(\cos\theta_{ij},\sin\theta_{ij})$ of the lattice site $j$ concerning to the site $i$, with $\delta_{i}$ and $\zeta_{j}$ respectively defined as $\eta_{i}=2\pi z_{i}/q$, $\zeta_{j}=2\pi\sigma_{j}/q$, with $z_{i}=0,\dots,q-1$, and $\sigma_{j}=0,\dots,q-1$. Each one of the real values $v_{m}$, $c_{p}$, and $f_{d}$ is related to a Reynolds’ rule; they correspond respectively to the velocity matching, collision avoidance, and flock centering parameters. For instance, setting $v_{m}>0$ we induce the bird to match the averaged observed state of its neighborhood; setting appropriate values for $c_{a}$ and $f_{d}$ we induce in the bird a tendency to penalize heading directions parallel and opposite to the averaged observed state ($\left|\eta_{i}-\zeta_{j}\right|=\pi$) both to avoid collisions when the verses of corresponding directions are convergent ($\theta_{ij}=\eta_{i}$), and to force the bird to stay in group when verses are divergent ($\theta_{ij}\neq\eta_{i}$). In our simulations, we used the following parameter values: $v_{m}=4$, $c_{a}=2$, and $f_{c}=1$.
@@
    where $\mathbf{G}(u_{t})$ is the expected free energy of the policy $u_{t}$ one-control-state long, and $\gamma\in\mathbb{R}$ is a variable denoted as “active inference precision”, on-line computed to self-tune the control-state selection process adaptively Starting from $\mathbf{G}$ definition as ELBO (Evidence Lower BOund) of the model evidence, it is possible to lead this expression back to already known distributions by using the position $Q(\sigma|z,u)\triangleq P(\sigma|z)$, where $Q$ denotes the variational approximation for the generative model. Through easy algebraic manipulations, we can write $\mathbf{G}$ as:
@@ A.3 A generative model variant for escaping from predator attacks
Individual vigilance, dilution and detection (foster1981evidence), the ”fountain effect” (hall1986predator), and confusion (neill1974experiments) are some of the escape strategies that rely on the benefits of being in a flock, where the probability of survival of any one individual increases when its behavior is configurable within a group response.
The choice of escape strategy depends on multiple factors (hilton1999intraflock), such as the predator’s approach— the closer the predator gets, the more likely the flock is to initiate an escape response—, the reaction time variation—the distance between predators and individuals influences their reaction time—, and the flock size, connected to predation risk of any individuals as highlithed also by the selfish herd theory (hamilton1971geometry).
Simulating a collective escape strategy entails extending the original flocking model: each bird must be able to recognize danger and modify behavior in safety situations. To this end, we introduce a second factor, named “stress state”, in the hidden state that encodes the presence of danger. Therefore, the hidden state of the extended model becomes $z^{i}=z^{i1}\otimes z^{i2}$, where $z^{i1}$ coincides with the hidden state of the original model, and $z^{i2}$ is a binary variable that indicates being ($z^{i2}=1$) or not ($z^{i2}=0$) in danger.
Analogously, we account for the observation of stress responses by adding, for each $\sigma^{j1}$, corresponding to the states of $j$-th neighbor individual, another factor encoding its stress outcome $\sigma^{j2}$, so that $\tilde{\sigma}=\tilde{\sigma}^{1}\otimes\tilde{\sigma}^{2}$, ultimately.
At this point, we need to characterize birds’ behavior in dangerous situations. We decided to implement the confusion strategy as an escape strategy. In practice, birds in the flock respond to a predator’s attack by fleeing at random. To do this, we need to modify the likelihood matrix $P(\mathbf{\tilde{\sigma}}|z^{i1},z^{i2})$ by extending it to the conditional probabilities of the observations $\tilde{\sigma}$ given the hidden state $z^{i2}$. When $z^{i2}=0$, the extended likelihood matrix is still defined as Equation (2), with $\sigma^{j}\equiv\sigma^{j1}$, while the threat-related component $\sigma^{j2}$ is uninformative (all the elements have the same value and every column sums to one). In contrast, when $z^{i2}=1$, the component $P(\sigma^{j1}|z^{i1},1)$ of the likelihood is uninformative; in contrast, the expression of $P(\sigma^{j2}|z^{i1},1)$ is analogous to that in Equation (2), taking care to replace the function $R(i,j)$ introduced in Equation (3) with the function $\bar{R}(i,j)$ that satisfies the following conditions:
@@
In a way, $z^{i2}$ is a context variable that shapes the perception of the states of neighboring birds and affects the choice of action. Actually, in the latter case, the influence is mutual. Indeed, similarly to the case without predators with $z_{t}^{i}\equiv z_{t}^{i1}$ (see Equation (4)), there exists a transition probability that probabilistically attributes a value to the stress state to $z_{t}^{i1}$, every time an action $u$ is executed.
Unlike $z_{t}^{i1}$, the transition model of $z_{t}^{i2}$ is conditioned by some parameters that entail how the stress state evolves. We presumed that the stress state could follow a specific dynamics that, from the initial threat, leads to a state of no danger present before the attack. Specifically, suppose to denote as $\bar{t}$ the time at which a generic bird has a predator (or a “stressed” bird) in its neighborhood. In that case, we hypothesize that its hidden state ${z}_{\bar{t}}^{i2}$ transitions from 0 to 1, then decays to 0 with a mean lifetime constant $\tau$. After that, the individual remains in a quiet state with $z_{t}^{i2}=0$ for a refractory period $T_{r}$, during which the individual does not respond to new danger situations. The following transition model can represent the entire dynamics over time:
@@
where the columns and rows of the matrices represent the values of $z_{t-1}^{i2}$ and $z_{t}^{i2}$, respectively. It should be noted that Equation (8) is independent of the action $u_{t}$ executed and, instead, depends exclusively on the time. In our simulations, we set $\tau=2$ and $T_{r}=10$.

### v2

@@ 5 Discussion
In the first analysis, we used the concept of (nested) Markov blankets to formalize how agency can extend beyond individual active inference agents to groups of agents—that is, from birds to flocks. We showed that during flocking, the Markov blanket, which formalizes a statistical boundary around an agent, can expand to enclose a set of birds forming a collective unit. Within the flock, subsets of birds play the roles of internal, sensory, and active states of a collective Markov blanket, while others constitute external states. This analysis extends prior work on collective, multi-agent systems—including groups of neurons (Palacios et al., 2019), collections of primordial cells (Friston, 2013), and developing tissues (Friston et al., 2015)—by showing that the notion of the Markov blanket can account for situations in which each agent retains its individual autonomy, while the collective simultaneously acquires a higher-level, joint agency. From this perspective, agency is not a fixed attribute of individual systems but an emergent property of coupled dynamics that maintain conditional independence from their surroundings. As interactions among agents strengthen, the boundaries that define “self” can expand or contract, giving rise to transient or stable forms of joint agency. This provides a unified statistical and dynamical framework for understanding how perception, action, and decision-making scale from individuals to collectives—from neurons forming neural assemblies to organisms forming social or ecological systems.
@@
The simulations revealed that larger flocks reacted more rapidly to the predator, as indicated by a faster onset of destabilization during later stages, even though the rules governing the propagation of “stress” did not depend on the level of organization. Most importantly, our results show that the flock as a whole possesses synergistic information about the predator’s location, and that this information increases with flock size and extends farther from the predator. This suggests a formally grounded—albeit implicit—notion of collective knowledge (or shared world model) that goes beyond the information available to individual birds.
@@
Summing up, our simulations illustrate a simple yet general method to characterize a collective agent, such as a flock, as an entity that maintains statistical separation from its environment through a Markov blanket, and exhibits knowledge beyond its components through synergistic information. More broadly, these results illustrate how collective entities can extend the spatial and temporal reach of perception, cognition, and action—or their cognitive light cone (Levin, 2019)—beyond the boundaries of any single agent.
Future work could extend and increase the realism of the flocking simulation in several ways. Although flocking in nature is a continuous process, here we model it as a discrete one. This choice follows a large body of work, which has shown that flocking dynamics can be accurately captured using discrete formulations based on lattice spin-glass models derived from statistical physics Solon and Tailleur (2013); Chatterjee et al. (2020). One advantage of this approach is that, while finite-size effects render numerical studies of continuous models computationally costly and analytical treatments generally complex Ginelli (2016), spin-glass formulations are more tractable while preserving key physical features of flocking behavior. In our setting, another key advantage is that the discrete formulation allows us to easily identify regions of aligned elements coexisting with regions of differing or disordered alignment, a distinction that is crucial for our assessment of nested Markov blankets. By contrast, the macroscopic properties of continuous models typically rely on a global transition from disordered to ordered phases, treated as a statistical property of the entire system Bialek et al. (2012). Nonetheless, future work could revisit our findings using continuous-time formulations of multi-agent active inference, as explored in previous studies (Friston, 2013; Friston et al., 2015; Heins et al., 2024). This would allow identification of Markov blankets by examining the spatial clusters that emerge and group agents, offering a complementary perspective to the methodology adopted here.
Future work could also revisit several simplifying assumptions of our simulation to increase biological realism. For instance, our model considers only flocking configurations in which each cell contains exactly one bird. While this assumption appropriately captures high-density regions of the flock—typically composed of interior individuals with uniformly distributed local interactions—it less accurately represents peripheral regions. Moreover, the model assumes a unit distance between neighboring cells. This simplifying assumption is motivated by evidence that interactions governing collective behavior follow a topological, rather than metric, framework Ballerini et al. (2008), whereby individuals interact with a fixed number of neighbors rather than maintaining fixed distances. Nevertheless, future work could explore ways to relax these assumptions in order to model more realistic flocking scenarios and to address novel questions, such as how sparseness and inter-individual distance influence the formation and dynamics of Markov blankets and synergistic information about predator location.
@@
Future studies could also address the important conceptual distinction between the implicit form of collective knowledge revealed by synergistic information and the explicit forms typically studied in cognitive science. In cognitive and social sciences, collective knowledge is often assumed to rely on explicit internal models of others’ mental states or shared epistemic representations (e.g., “common ground” or “what we all know that we know”). Similarly, our simulations demonstrate collective sensing and action that do not depend on such explicit representations but instead arise from simple interaction rules. In human societies, collective action is often defined by intentionality—individuals act with the goal of producing collective outcomes. By contrast, in simpler biological systems, such as morphogenesis or pattern formation, intentionality can be replaced by teleological organization: systems are structured in such a way to produce collective outcomes even without explicit goals.
@@
An open question, then, is how to bridge our minimal notion of joint agency with more sophisticated accounts used in cognitive science. A natural starting point is the good regulator theorem (Conant and Ross Ashby, 1970), which states that any effective controller must embody a model of the system it regulates. In the multi-agent flock we describe, no single bird possesses an explicit generative model of the flock’s collective behavior or its interaction with the environment. Instead, the flock’s generative model is implicitly distributed across the birds and their interaction patterns. In this sense, one can interpret this as the flock not having a model of collective action but rather being—or dynamically becoming—such a model through the structured couplings among its constituent agents. More complex agents, by contrast, possess explicit internal models of collective behavior, with shared representations, goals, and intentions. Understanding how living systems evolve from being to having a model may shed light on the developmental and evolutionary origins of collective cognition (Pezzulo et al., 2022).
Another possible line of work could explore potential cross-fertilizations between the simple simulations presented here and recent efforts to understand more complex forms of collective cognition from an inferential perspective. For example, Collective Predictive Coding investigates how agents form communities and develop shared linguistic symbols through decentralized, independent action decisions (Taniguchi et al., 2025a, b; Taniguchi, 2024). Another perspective, Thinking Through Other Minds, emphasizes the emergence of shared understanding and social coordination by modeling how agents represent and infer the mental states of others (Veissière et al., 2020). Yet another approach examines long-lasting societal dynamics within groups of agents, using large language models as generative models (Park et al., 2023). What distinguishes our models from these approaches is the relative simplicity of the agents’ generative models and their (non-linguistic) forms of communication. However, in principle, the common framework of multi-agent generative modeling could allow the same methodology to be applied across a wide range of collective behaviors. Future work could investigate whether and how the methods used here to characterize the emergence of joint agency, (implicit or explicit) collective knowledge, and shared world models in simple multi-agent systems, such as flocks, can be extended to study more sophisticated forms of collective cognition.
@@
Finally, this perspective invites consideration of agency at even broader spatial and temporal scales, such as in niche construction (Constant et al., 2018; Pio-Lopez et al., 2025). Extending these ideas across generations and cultural dynamics raises deeper questions (Pezzulo et al., 2025b; Clark and Chalmers, 1998; Friston et al., 2024): To what extent can the nesting of Markov blankets help us understand interactions between individual and social cognition across scales, the social and cultural dynamics of human societies, the formation of collective agency and extended minds, and the ways in which knowledge—distributed across people, artifacts, and institutions such as books, tools, and the internet—shapes individual cognition? We leave these questions open for future research.
@@ A.1 Active inference framework
In this work, we consider birds as Bayesian agents with sensory, active, and internal states that are updated according to the Active Inference principles (Parr et al., 2022). Active inference (AIF) is a theoretical framework that provides a unified account of perception, action, and learning of both living and artificial systems. It posits that an agent’s behavior can be understood as the process of maximizing the evidence for the implicit statistical model of the world it embodies, in response to the sensory information flow, by selecting adaptive sequences of actions.
@@
We adopted a hybrid modeling approach. We borrowed the Potts model (Wu, 1982), a statistical physics model describing spin glasses in which spins are arranged on a lattice, a periodic graph, and combined it with the computational model developed by Reynolds to mimic collective behaviors in computer vision applications (Reynolds, 1987).
The Potts model is a generalization of the Ising model (Ising, 1925) for $q>2$ components. It has been used to address numerous problems in collective behavior, which are often described as instances of lattice statistics. Indeed, replacing continuous symmetry with discrete symmetry allows for a simpler, more tractable understanding of the flocking transition. Recently, Solon and Tailleur argued that flocking can be described as a transition from a disordered phase to a polar-ordered phase in an active Ising model, in which spins both diffuse and align on the lattice, a coarse-grained representation of the space (Solon and Tailleur, 2013). (Chatterjee et al., 2020) studied a square lattice in which active particles have four internal states corresponding to the four directions of motion. A local alignment rule inspired by the ferromagnetic 4-state Potts model, combined with self-propulsion via biased diffusion based on the internal particle states, leads to flocking at high densities and low noise.
@@ A.2 Generative model for flocking
The energy of a system represented in this way coincides with the Hamiltonian of a Potts model on a lattice spin glass, defined as
$\mathcal{H}=J_{P}\sum_{(i,j)}\delta(z_{i},z_{j})$ (1)
where $\delta(z_{i},z_{j})$ is the Kronecker delta, which equals one whenever $z_{i}=z_{j}$ and zero otherwise, $(i,j)$ are the sets of the $M$ nearest neighbor pairs for the indexes $i$, and $J_{P}$ is a coupling constant that depends on $q$, and is equal for all the potential configurations.
@@
    where $\theta$ is the angle associated to the two-dimensional unit position vector $(\mathbf{r}_{j}-\mathbf{r}_{i})=(\cos\theta,\sin\theta)$, and where $\eta_{i}$ and $\zeta_{j}$ are respectively defined as $\eta_{i}=2\pi z_{i}/q$, $\zeta_{j}=2\pi\sigma_{j}/q$, with $z_{i}=0,\dots,q-1$, and $\sigma_{j}=0,\dots,q-1$. Each one of the real values $v_{m}$, $c_{a}$, and $f_{c}$ is related to a Reynolds’ rule; they correspond respectively to the velocity matching, collision avoidance, and flock centering parameters. For instance, setting $v_{m}>0$ we induce the bird to match the averaged observed state of its neighborhood; setting appropriate values for $c_{a}$ and $f_{c}$ we induce in the bird a tendency to penalize heading directions parallel and opposite to the averaged observed state ($\left|\eta_{i}-\zeta_{j}\right|=\pi$) both to avoid collisions when the verses of corresponding directions are convergent ($\theta=\eta_{i}$), and to force the bird to stay in group when verses are divergent ($\theta_{ij}\neq\eta_{i}$). In our simulations, we used the following parameter values: $v_{m}=4$, $c_{a}=2$, and $f_{c}=1$.
@@
    where $\mathbf{G}(u_{t})$ is the expected free energy of the policy $u_{t}$ one-control-state long, and $\gamma\in\mathbb{R}$ is a variable denoted as “active inference precision”, on-line computed to self-tune the control-state selection process adaptively. Starting from $\mathbf{G}$ definition as ELBO (Evidence Lower BOund) of the model evidence, it is possible to lead this expression back to already known distributions by using the position $Q(\sigma|z,u)\triangleq P(\sigma|z)$, where $Q$ denotes the variational approximation for the generative model. Through easy algebraic manipulations, we can write $\mathbf{G}$ as:
@@ A.3 A generative model variant for escaping from predator attacks
Individual vigilance, dilution and detection (Foster and Treherne, 1981), the ”fountain effect” (Hall et al., 1986), and confusion (Neill and Cullen, 1974) are some of the escape strategies that rely on the benefits of being in a flock, where the probability of survival of any one individual increases when its behavior is configurable within a group response.
The choice of escape strategy depends on multiple factors (Hilton et al., 1999), such as the predator’s approach— the closer the predator gets, the more likely the flock is to initiate an escape response—, the reaction time variation—the distance between predators and individuals influences their reaction time—, and the flock size, connected to predation risk of any individuals as highlithed also by the selfish herd theory (Hamilton, 1971).
Simulating a collective escape strategy entails extending the original flocking model: each bird must be able to recognize danger and modify behavior in safety situations. To this end, we introduce a second factor, named “stress state”, in the hidden state that encodes the presence of danger. Therefore, the hidden state of the extended model becomes $z^{\prime i}=z^{i}\otimes z^{*i}$, where $z^{i}$ coincides with the hidden state of the original model, and $z^{*i}$ is a binary variable that indicates being ($z^{*i}=1$) or not ($z^{*i}=0$) in danger.
Analogously, we account for the observation of stress responses by adding, for each $\sigma^{j}$, corresponding to the states of $j$-th neighbor individual, another factor encoding its stress outcome $\sigma^{*j}$, so that $\sigma^{\prime j}=\sigma^{j}\otimes\sigma^{*j}$, ultimately.
At this point, we need to characterize birds’ behavior in dangerous situations. We decided to implement the confusion strategy as an escape strategy. In practice, birds in the flock respond to a predator’s attack by fleeing at random. To do this, we need to modify the likelihood matrix $P(\mathbf{\tilde{\sigma}}|z^{i},z^{*i})$ by extending it to the conditional probabilities of the observations $\tilde{\sigma}$ given the hidden state $z^{*i}$. When $z^{*i}=0$, the extended likelihood matrix is still defined as Equation (3), with $\sigma^{\prime j}\equiv\sigma^{j}$, while the stress component $\sigma^{*j}$ is uninformative (all the elements have the same value and every column sums to one). In contrast, when $z^{*i}=1$, the component $P(\sigma^{\prime j}|z^{i},1)$ is analogous to that in Equation (3), taking care to replace the function $R(i,j)$ introduced in Equation (4) with the function $R^{\prime}(i,j)$ that satisfies the following conditions:
@@
In a way, $z^{*i}$ is a context variable that shapes the perception of the states of neighboring birds and affects the choice of action. Actually, in the latter case, the influence is mutual. Indeed, similarly to the case without predators with $z_{t}^{\prime i}\equiv z_{t}^{i}$ (see Equation (5)), there exists a transition probability that probabilistically attributes a value to the stress state to $z_{t}^{i}$, every time an action $u$ is executed.
Unlike $z_{t}^{i}$, the transition model of $z_{t}^{*i}$ is conditioned by some parameters that entail how the stress state evolves. We presumed that the stress state could follow a specific dynamics that, from the initial threat, leads to a state of no danger present before the attack. Specifically, suppose to denote as $\bar{t}$ the time at which a generic bird has a predator (or a “stressed” bird) in its neighborhood. In that case, we hypothesize that its hidden state ${z}_{\bar{t}}^{*i}$ transitions from 0 to 1, then decays to 0 with a mean lifetime constant $\tau$. After that, the individual remains in a quiet state with $z_{t}^{*i}=0$ for a refractory period $T_{r}$, during which the individual does not respond to new danger situations. The following transition model can represent the entire dynamics over time:
@@
where the columns and rows of the matrices represent the values of $z_{t-1}^{*i}$ and $z_{t}^{*i}$, respectively. It should be noted that Equation (9) is independent of the action $u_{t}$ executed and, instead, depends exclusively on the time. In our simulations, we set $\tau=2$ and $T_{r}=10$.